# Towards Urban Planing AI Agent in the Age of Agentic AI

### v2

@@ I Introduction and Motivation
Urban planning is an interdisciplinary and complex process that involves with public policy, social science, engineering, architecture, landscape, and other related fields. Effective urban planning can help to mitigate the operational and social vulnerability of a urban system, such as high tax, crimes, traffic congestion and accidents, pollution, depression, and anxiety.
@@
Recent breakthroughs in Generative AI (GenAI), Large Language Models (LLMs), and Agentic AI offer a transformative opportunity to rethink urban planning. These technologies introduce the abilities to generate alternative futures, understand complex textual and visual data, and act as autonomous agents that reason, simulate, and navigate planning goals. The new abilities of these AI paradigms makes it possible to develop an AI Urban Planner that can augment human expertise, democratize access to planning insights, and adapt to changing urban conditions. This insight prompts a fundamental question: how might AI reshape the practice and ethics of urban planning in the era of AI? What roles does GenAI, LLMs, and agentic AI play in urban planning? Can machines develop and learn at a human capability to automatically and quickly calculate land-use configurations? In this way, machines can be planning assistants and human planners can adjust machine-generated plans for specific needs.
@@ II Existing Studies: Urban Planning as Generative AI
We propose to leverage and adapt these GenAI models to produce diverse land-use configurations and generate spatial designs conditioned on multimodal inputs, such as urban geography data, human mobility patterns, and environmental constraints. One example task is to create alternative zoning proposals that meet goals for density, connectivity, or climate resilience, while incorporating learned priors from historical planning outcomes. Such new capabilities enable urban planners to shift from rule-based prescriptions to data-driven generative design that is both adaptive and customizable.
@@ II-B2 Generation: Producing Optimal Land-Use and Built Environment Configurations
Given a learned representation, the generative stage synthesizes land-use plans using deep generative models, such as, Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), Autoregressive Models, Flow-Based Models, Energy-Based Models. These generatie urban planning models can be trained with a combination of generic objectives (e.g., reconstruction loss, likelihood maximization, adversarial loss) and domain-specific constraints tailored to urban planning. These domain-specific constraints balance a diverse mixture of objectives spanning spatial, social, economic, environmental, and governance domains: 1) Spatial objectives focus on land-use compatibility, connectivity, and density optimization to support functional and accessible urban form. 2 )Social objectives emphasize equity, inclusion, community cohesion, and housing affordability to foster livable and just environments. accessible urban form. 3) Economic objectives include maximizing land value, supporting local economies, and minimizing infrastructure costs. 4) Environmental objectives aim to enhance sustainability, reduce emissions, and preserve natural ecosystems. 5) Governance objectives demand compliance with legal frameworks, responsiveness to stakeholder input, and adaptability over time. These objectives often conflict, for example, between maximizing density and preserving green space—highlighting the need for AI systems that can model, navigate, and transparently negotiate such multi-objective trade-offs in planning support tools.
Together, representation and generation form a pipeline where AI systems learn to ”understand” the urban environment and ”create” planning proposals that are not only technically valid but socially and contextually meaningful.
@@ II-D Limitations of Current AI Urban Planner Research
Despite advances in AI for urban planning, current approaches face critical limitations across five foundational dimensions: theory, methodology, data, computing, and applications. 1) Theory-wise, most AI models reduce planning to spatial optimization or land-use generation, thus, ignore core urban concepts such as zoning, resilience, social equity, and participatory governance, and often ignore the political and normative nature of planning decisions—undermining interpretability and legitimacy. 2) Method-wise, models [2, 6, 5, 4, 1, 3, 2] rely on single-step or loosely staged generation conditioned on static inputs that limit creativity and adaptivity, fail to capture multi-granularity dynamics, multiple planning objective trade-offs, and adaptive processes. 3) Data-wise, existing city-specific datasets (e.g., Beijing, NYC) lacks clear evaluation criteria on scoring a good planning and a bad planning. 3) Computation-wise, GenAI models are resource-intensive and prone to instability, especially in adversarial or variational settings; they lacks explainability and transparency due to the use of neural network based architectures, and often remain opaque to end users. 4) Application-wise, existing generative planning methods ignore many domain expert developed tools. Be sure to notice that domain urban planners have developed various tools to assist each step of urban planning process guided by planning theory, while existing pure neural networks based generation ignore the power of the tools developed by urban planner practitioners.
@@ III The Vision: Integrating VLM, Multimodal Reasoning, and Agentic AI for Urban Planning AI Agents
To go beyond GenAI like VAEs, GANs, or diffusion models, we believe that VLMs integrate visual and textual modalities and enable more semantically grounded, instruction-following generation. For example, we can align urban design prompts (e.g., “generate a transit-oriented mixed-use layout”) with satellite imagery, GIS data, and street-level visuals, in order to allow planners to interact with models using natural language while preserving spatial fidelity. Agentic AI further models the complex urban planning pipeline by introducing planning, reasoning, task decomposition, and multi-step decision-making abilities. Agent-based urban planners can autonomously explore spatial trade-offs, simulate long-term effects, and iteratively refine plans based on feedback, environmental constraints, or human instructions. VLMs and agentic systems enable an AI urban planner that is not only generative, but goal-driven, context-aware, and aligned with multimodal planning tasks that are capable of understanding complex instructions, grounding outputs in real-world data, and dynamically adapting to evolving objectives.
[To BE FINISHED]

### v3

@@ I Introduction and Motivation
Urban planning is an interdisciplinary and complex process that involves with public policy, social science, engineering, architecture, landscape, and other related fields. Effective urban planning can help to mitigate the operational and social vulnerability of a urban system, such as high tax, crimes, traffic congestion and accidents, pollution, depression, and anxiety [1, 2, 3].
@@
Recent breakthroughs in Generative AI (GenAI) [4, 5, 6, 7], Large Language Models (LLMs) [8, 9, 10], and Agentic AI [11, 12] offer a transformative opportunity to rethink urban planning. These technologies introduce the abilities to generate alternative futures, understand complex textual and visual data, and act as autonomous agents that reason, simulate, and navigate planning goals. The new abilities of these AI paradigms makes it possible to develop an AI Urban Planner that can augment human expertise, democratize access to planning insights, and adapt to changing urban conditions. This insight prompts a fundamental question: how might AI reshape the practice and ethics of urban planning in the era of AI? What roles does GenAI, LLMs, and agentic AI play in urban planning? Can machines develop and learn at a human capability to automatically and quickly calculate land-use configurations? In this way, machines can be planning assistants and human planners can adjust machine-generated plans for specific needs.
@@ II Existing Studies: Urban Planning as Generative AI
We propose to leverage and adapt these GenAI models to produce diverse land-use configurations and generate spatial designs conditioned on multimodal inputs, such as urban geography data, human mobility patterns, and environmental constraints [13]. One example task is to create alternative zoning proposals that meet goals for density, connectivity, or climate resilience, while incorporating learned priors from historical planning outcomes. Such new capabilities enable urban planners to shift from rule-based prescriptions to data-driven generative design that is both adaptive and customizable [14].
@@ II-B2 Generation: Producing Optimal Land-Use and Built Environment Configurations
Given a learned representation, the generative stage synthesizes land-use plans using deep generative models, such as, Variational Autoencoders (VAEs) [21], Generative Adversarial Networks (GANs) [22, 23, 24], Autoregressive Models [25, 26], Flow-Based Models, Energy-Based Models. These generative urban planning models can be trained with a combination of machine learning losses (e.g., reconstruction loss, likelihood maximization, adversarial loss) and application-specific regularization tailored to urban planning. Together, representation and generation form a pipeline where AI systems learn to ”understand” the urban environment and ”create” planning proposals that are not only technically valid but socially and contextually meaningful.
@@ II-D Limitations of Current AI Urban Planner Research
Despite advances in AI for urban planning, current approaches face critical limitations across five foundational dimensions: theory, methodology, data, computing, and applications. 1) Theory-wise, most AI models reduce planning to spatial optimization or land-use generation, thus, ignore core urban concepts such as zoning, resilience, social equity, and participatory governance, and often ignore the political and normative nature of planning decisions—undermining interpretability and legitimacy. 2) Method-wise, models [16, 20, 19, 18, 15, 17, 28] rely on single-step or loosely staged generation conditioned on static inputs that limit creativity and adaptivity, fail to capture multi-granularity dynamics, multiple planning objective trade-offs, and adaptive processes. 3) Data-wise, existing city-specific datasets (e.g., Beijing [29], NYC) lacks clear evaluation criteria on scoring a good planning and a bad planning. 3) Computation-wise, GenAI models are resource-intensive and prone to instability, especially in adversarial or variational settings; they lacks explainability and transparency due to the use of neural network based architectures, and often remain opaque to end users. 4) Application-wise, existing generative planning methods ignore many domain expert developed tools. Be sure to notice that domain urban planners have developed various tools to assist each step of urban planning process guided by planning theory, while existing pure neural networks based generation ignore the power of the tools developed by urban planner practitioners.
@@ III The Vision: Integrating VLM, Multimodal Reasoning, and Agentic AI for Urban Planning AI Agents
To go beyond GenAI like VAEs, GANs, or diffusion models, we believe that VLMs integrate visual and textual modalities and enable more semantically grounded, instruction-following generation [30, 31]. For example, we can align urban design prompts (e.g., “generate a transit-oriented mixed-use layout”) with satellite imagery, GIS data, and street-level visuals, in order to allow planners to interact with models using natural language while preserving spatial fidelity. Agentic AI further models the complex urban planning pipeline by introducing planning, reasoning, task decomposition, and multi-step decision-making abilities. Agent-based urban planners can autonomously explore spatial trade-offs, simulate long-term effects, and iteratively refine plans based on feedback, environmental constraints, or human instructions. VLMs and agentic systems enable an AI urban planner that is not only generative, but goal-driven, context-aware, and aligned with multimodal planning tasks that are capable of understanding complex instructions, grounding outputs in real-world data, and dynamically adapting to evolving objectives.
Despite the interdisciplinarity and complex nature of urban planning, arising from its multifaceted requirements, agentic AI introduces a new paradigm that directly responds to these challenges. The integration of reasoning, task decomposition, and domain-specific tool usage empowers agentic AI to revolutionize urban planning towards more transparent, explainable, and practical urban configurations.
Reasoning plays a crucial role in this paradigm, enabling agentic AI to go beyond static generation.By systematically analyzing the diverse mixture of planning objectives, constraints, and potential conflicts, an AI urban planner can apply logical and context-aware reasoning to explicitly balance trade-offs across multiple urban planning requirements, thereby facilitating informed decision-making, anticipating potential downstream consequences, and producing planning outcomes that are both technically feasible and contextually meaningful.
Task decomposition enables agentic AI to effectively address the inherent complexity of urban planning. Instead of generating a comprehensive urban planning alternative in a single step, agentic AI system adopts a divide-and-conquer paradigm that decomposes the planning process into manageable sub-tasks, where each sub-task focuses on a specific dimension of the planning. Importantly, each sub-task can also be governed by its own objectives and constraints, allowing the system to incorporate local requirements while maintaining consistency at the system level, ensuring more transparent and traceable intermediate decisions.
Domain-specific tool usage further enhances agentic AI by integrating practical tools developed by urban planning practitioners. These tools have been repeatedly validated in real-world projects and regulatory processes, including decades of planning theory and professional practice. By leveraging such domain-expert designed tools, agentic AI system can align the outputs with up-to-date institutional standards and long-standing urban planning principles rather than relying solely on static data-driven inference, facilitating innovative, credible and practically applicable urban planning alternatives.These capabilities form a comprehensive pipeline that effectively analyzes, summarizes, and generates optimal planning decisions under evolving urban objectives, thereby addressing the inherent complexity of urban planning and rendering agentic AI systems more practical and advanced.
Figure 1: Framework Overview of Agentic AI based Automated Urban Planner.
Building on these capabilities, we conceptualize an agentic AI urban planner paradigm. Figure 1 demonstrates the overall pipeline of an agentic AI urban planner. 1) Task Planner Agent: Given the heterogeneous objectives and constraints, the agentic AI system starts with a Task Planner agent. The task planner agent has three primary objectives. First, it develops a comprehensive understanding of the overall planning task by integrating the diverse requirements and constraints. Since the requirements are highly heterogeneous, the input to the task planner agent might have different forms such as textual documents, images, spatial data, and numerical indicators. This diverse nature necessitates the capability of interpreting and incorporating multimodal inputs. Multimodal abstraction is essential for capturing the dependencies between objectives, constraints, and contextual factors. Second, it decomposes the intricate planning task by adopting a divide-and-conquer paradigm. Instead of generating the planning alternative as a whole in a single step, the task planner agent partitions the process into a series of smaller and more manageable sub-tasks. Each task represents a specific dimension of the planning task and is associated with its own objectives and local constraints, while remaining consistent with the overarching global goals. Third, it makes well-informed decisions on sub-tasks and sub-constraints assignment. The task planner agent determines how each sub-task should be delegated to specialized agents to ensure that the assigned task matches the expertise, computational capabilities, and available domain-specific tools of specialized agents. For instance, given the previous example of redevelopment zone along the Gulf Coast where agentic AI urban planner reconfigures urban layouts toward improved flood resilience in a divide-and-conquer manner, the task planner agent first integrates multimodal inputs, including flood simulations, infrastructure risk maps, and planner-defined intents, to form a comprehensive understanding of the planning context. Based on the understanding, the task planner agent decomposes the multi-objective planning task into a set of tractable sub-tasks, including flood risk modeling, critical infrastructure exposure analysis, NbS scheme generation, and cost–benefit analysis. 2) Specialized Agents: Once the planning task has been decomposed, the resulting sub-tasks are assigned to Specialized Agents for execution. Each specialized agent is dedicated to a clearly defined scope of work, focusing on one particular dimension of the planning task. Instead of addressing the full complexity of the planning task, they operate within the narrower boundaries of their designated sub-tasks, enabling more accurate sub-outputs. Moreover, the specialized agents are empowered with domain-specific methods and practitioner-developed tools so that they have the access to professional expertise and regulatory standards. For instance, in the previous example, it requires 4 specialized agents:
  • Hydrology Agent applies hydrological and hydraulic models (e.g., HEC-RAS, SWMM) to generate flood inundation maps across different return periods
  • Infrastructure Agent employs GIS overlays and facility databases to evaluate the exposure of hospitals, schools, and power stations
  • NbS Agent leverages green infrastructure siting models to identify potential wetland restoration zones, rain garden networks, and green-blue corridors
  • Economics Agent utilizes cost–benefit analysis models to evaluate trade-offs between flood damage reduction and investment cost for each alternative.
3) Fusion Agent: After the specialized agents have completed their corresponding sub-tasks, their outputs are fed forward to the fusion agent. The fusion agent then aggregates the sub-outputs from the specialized agents to avoid potentially conflicting sub-results and synthesize them into a coherent intermediate plan, ensuring internal consistency across spatial, social, and environmental dimensions. By integrating these results, it produces planning alternatives that balance flood protection, infrastructure security, ecological integrity, and financial feasibility. In this way, the fusion agent serves as an intermediate system-level integrator, transforming fragmented sub-outputs into consistent and feasible planning candidates. This agent is critical because specialized agents, while accurate within their own specialized scope, may produce outputs that are misaligned. For instance, an economically optimal solution may conflict with environmental resilience or social equity requirements. In the flooding resilience case, the fusion agent combines flood inundation maps from the hydrology agent, vulnerability assessments from the infrastructure agent, ecological restoration options from the NbS agent and cost–benefit outcomes from the economics agent. 4) Multimodal Generation Agent: Once a set of coherent planning candidates has been synthesized by the fusion agent, a multimodal generation agent eventually transforms the planning candidates into multimodal outputs that are transparent, interpretable and accessible. Instead of solely reporting textual contexts, the multimodal generation agent can produce complementary formats such as narrative summaries, tabular statistics, spatial maps and model configurations, ensuring that both technical experts and non-specialist audiences can effectively engage with the results. For example, in the flooding resilience case, the multimodal generation agent may generate visual flood risk and land-use maps comparing current and improved scenarios, metrics tables highlighting the exposure of critical infrastructure and textual reports comparing alternative NbS schemes (e.g., wetlands vs. rain gardens). In contrast to conventional automated urban planning approaches that often formulate urban planning as a static, one-shot optimization or generation task, the agentic AI paradigm reconceptualizes urban planning as a multi-step, reasoning-driven process. 1) Agentic AI incorporates reasoning and task decomposition to manage the inherent complexity of planning, while previous approaches rely on static and loosely staged generation that fail to capture multi-granularity dynamics and adaptive trade-offs. 2) Agentic AI empowers the specialized agents with domain-specific methods and practitioner-developed tools, ensuring that outputs align with regulatory standards and professional expertise, where traditional models largely ignore decades of planning theory and expert practices. 3) Agentic AI enhances transparency and interpretability by coordinating sub-tasks and integrating them through a fusion agent, while prior approaches often formulate planning as a black-box optimization that obscures trade-offs and undermines legitimacy 4) Agentic AI produces multimodal outputs such as textual reports, tabular summaries, and spatial maps, demonstrating planning alternatives through comprehensive representations and ensuring interpretability, systematic assessment, and practical utility for both experts and the public, whereas traditional approaches typically only yield abstract reconfigurations such as spatial layouts or land-use maps that are mathematically optimized and fails to express the implications of planning decisions from multiple dimensions or perspectives.